OpenStudy (anonymous):

i need two answers for cosθ + √3 = 0 if 0≤θ<300

OpenStudy (anonymous):

No calculators?

OpenStudy (anonymous):

calculators allowed

hartnn (hartnn):

then subtract \(\sqrt{3}\) from both sides to get \(cos \theta=-\sqrt{3} \) use calculator to find theta.

OpenStudy (anonymous):

i did, but it says Math Error

hartnn (hartnn):

then please verify the question: it must be 2cos theta + root 3

OpenStudy (anonymous):

correct